# Variational Flow Matching for Graph Generation

### v1

@@ 4 Flow Matching and Score Matching: Bridging the Gap
Informally, the following theorem states that VFM has flow matching as a special case when the variational approximation is Gaussian and under certain assumptions on $u_{t}(x\mid x_{1})$. This result is key, as it shows that indeed optimizing a model through VFM is more flexible than regular flow matching. Formally, the following holds (see theorem 3 in section A.2 for the proof):
@@
Flow matching [24] is inspired by score-based models [46] and shares strong connections with them. In score-based models, the objective is to approximate the score function $\nabla_{x}\log p_{t}(x)$ with a function $s^{\theta}_{t}(x)$. A connection to VFM becomes apparent by observing that the score function can also be expressed as an expectation with respect to $p_{t}(x_{1}\mid x)$ (see section A.5 for derivation):
@@
It is now clear that $s_{t}^{\theta}(x)=\nabla_{x}\log p_{t}(x)$ when $q_{t}^{\theta}(x_{1}\mid x)=p_{t}(x_{1}\mid x)$. This suggests that there exists a variational formulation of score-based models that is entirely analogous to VFM. More specifically, we see that both models could be optimized using a single objective, which yields an approximation $q_{t}^{\theta}(x_{1}\mid x)$ that parameterizes both $v_{t}^{\theta}(x)$ and $s^{\theta}_{t}(x)$.
Following [46, 3], we can construct stochastic generative dynamics $dx=\tilde{v}_{t}^{\theta}(x)dt+g_{t}dw$ to approximate the true dynamics $dx=\tilde{u}_{t}(x)dt+g_{t}dw$ (see details in section A.6), with
@@
This connection has two important implications. First, it shows that learning a variational approximation $q^{\theta}_{t}(x_{1}\mid x)$ can be used to define both deterministic and stochastic dynamics, whereas flow matching typically considers deterministic dynamics only (as flows are viewed through the lens of ODEs). Second, it enables us to show that a reweighted version of the VFM objective provides a bound on the log-likelihood of the model. This bound provides another theoretical motivation for learning using the VFM objective.
@@
We provide a proof in section A.7. We further note that that applying the same linearity condition that we discussed in section 3.2 to the conditional score function maintains all the same connections with score-based models.
@@ 5 Related Work
Several approaches to discrete generation using diffusion models have been developed [4, 28, 53, 17]. For graph generation specifically, [50] utilize a Markov process that progressively edits graphs by adding or removing edges and altering node or edge categories and is trained using a graph transformer network, that reverses this process to predict the original graph structure from its noisy version. This approach breaks down the complex task of graph distribution learning into simpler node and edge classification tasks. Moreover, [19] proposes a score-based generative model for graph generation using a system of stochastic differential equations (SDEs). The model effectively captures the complex dependencies between graph nodes and edges by diffusing both node features and adjacency matrices through continuous-time processes. These are non-autoregressive graph generation approaches that perform on par with autoregressive ones, such as in [26, 23, 31]. Other non-autoregressive approaches worth mentioning are [25, 29, 28].
@@
Very recently two flow-based methods for discrete generative modeling have been proposed, which differ both in terms of technical approach and intended use case from the work that we present here. [^2] Flow matching and diffusion models have also been proposed for geometric graph generation, e.g. in [20, 47] and [16, 49] respectively, but since these approaches are continuous (as they generate coordinates based on some conformer) they consider a fundamentally different task than the ones we consider here.
In [48], a Dirichlet flow framework for DNA sequence design is introduced, utilizing a transport problem defined over the probability simplex. This approach differs from CatFlow in that it represents the conditional probability path $p_{t}(x\mid x_{1})$ using a Dirichlet distribution. This implies that points $x$ are constrained to the simplex, which is not the case for CatFlow. Dirichlet Flows have not been evaluated on graph generation, but we did carry out preliminary experiments based on released source code. We have opted not to report these results; we did not obtain good performance out of the box, but have also not invested substantial time in architecture and hyperparameter selection.
In [6], Discrete Flow Models (DFMs) are introduced. DFMs use Continuous-Time Markov Chains to enable flexible and dynamic sampling in multimodal generative modeling of both continuous and discrete data. Though sharing a goal, this approach differs significantly from CatFlow as in the end the resulting model does not learn a CNF, but rather generation through sequential sampling from a time-dependent categorical distribution. As in the case of Dirichlet flows, no evaluation on graph generation was performed.
The switch to the variational perspective is inspired by [50], showing significant improvement through viewing the dynamics as a classification task over end points. However, CatFlow is still a continuous model, and integrates – rather than iteratively samples – during generation.
@@
GraphVAE 0.130 0.170 0.050 0.350 0.980 0.540
GNF 0.030 0.100 0.001 0.200 0.200 0.110
EDP-GNN 0.052 0.093 0.007 0.053 0.144 0.026
GDSS 0.021 0.024 0.007 0.045 0.086 0.007
@@ 6 Experiments
We evaluate CatFlow in three sets of experiments. First, we consider an abstract graph generation task proposed in [30], where the goal of this task is to evaluate if CatFlow is able to capture the topological properties of graphs. Second, we consider two common molecular benchmarks, QM9 [34] and ZINC250k [18], consisting of small and (relatively) large molecules respectively. This task is chosen to see if CatFlow can learn semantic information in graph generation, such as molecular properties. Finally, we perform an ablation comparing CatFlow to standard flow matching, specifically in terms of generalization. The experimental setup and model choices are provided in appendix C.
@@
MoFlow 91.36 98.65   4.467 63.11 99.99 20.931
EDP-GNN 47.52 99.25   2.680 82.97 99.79 16.737
GraphEBM   8.22 97.90   6.143   5.29 98.79 35.471
GDSS 95.72 98.46   2.900 97.01 99.64 14.656
Digress 99.00 96.20 - - - -
@@ 6.1 Abstract graph generation
We first evaluate CatFlow on an abstract graph generation task, including synthetic and real-world graphs. We consider 1) Ego-small (200 graphs), consisting of small ego graphs drawn from a larger Citeseer network dataset [41], 2) Community-small (100 graphs), consisting of randomly generated community graphs, 3) Enzymes (587 graphs), consisting of protein graphs representing tertiary structures of the enzymes from [40], and 4) Grid (100 graphs), consisting of 2D grid graphs. We follow the standard experimental setup popularized by [54] and hence report the maximum mean discrepancy (MMD) to compare the distributions of degree, clustering coefficient, and the number of occurrences of orbits with 4 nodes between generated graphs and a test set. Following [19], we also use the Gaussian Earth Mover’s Distance kernel to compute the MMDs instead of the total variation.
The results of the Ego-small and Community-small tasks are summarized in table 1, additional results (and error bars) are provided in appendix B. The results indicate that CatFlow is able to capture topological properties of graphs, and performs well on abstract graph generation.
@@ 6.2 Molecular Generation: QM9 & ZINC250k
Molecular generation entails designing novel molecules with specific properties, a complex task hindered by the vast chemical space and long-range dependencies in molecular structures. We evaluate CatFlow on two popular molecular generation benchmarks: QM9 and ZINC250k [34, 18].
We follow the standard setup – e.g. as in [42, 28, 50, 19] – of kekulizing the molecules using RDKit [22] and removing the hydrogen atoms. We sample 10,000 molecules and evaluate them on validity, uniqueness, and Fréchet ChemNet Distance (FCD) – evaluating the distance between data and generated molecules using the activations of ChemNet [33]. Here, validity is computed without valency correction or edge resampling, hence following [55] rather than [42, 28], as is more reasonable due to the existence of formal charges in the data itself. We do not report novelty for QM9 and ZINC250k, as QM9 is an exhaustive list of all small molecules under some chemical constraint and all models obtain (close to) 100% novelty on ZINC250k. [^3] CatfFlow obtains 49% novelty on QM9.
The results are summarized in table 2 and samples from the model are shown in fig. 2. CatFlow obtains state-of-the-art results on both QM9 and ZINC250k, virtually obtaining perfect performance on both datasets. It is worth noting that CatFlow also converges faster than flow matching and is not computationally more expensive than any of the baselines either during training or generation.
@@ 6.3 CatFlow Ablations
Figure 3: Ablation results, comparison between standard flow matching and CatFlow.
@@
In fig. 3 we report a score, which is the percentage of generated molecules that is valid and unique. CatFlow not only outperforms regular flow matching in the large model and full data setting, but is also significantly more robust to a decrease in model-size and data. Moreover, we observe significantly faster convergence (curves not shown). We hypothesize this is a consequence of the optimization procedure not exploring ‘irrelevant’ paths that do not point towards the probability simplex.
@@ A.1 Derivation of the Variational Flow Matching Objective
We derive eq. 9 that states the equivalence of the optimization of the VFM objective and minimization of the KL divergence between the true endpoint distribution $p_{t}(x_{1}\mid x)$ and the variational approximation $q^{\theta}_{t}(x_{1}\mid x)$. Note that
@@
Therefore, the second part of eq. 27 corresponds to negative cross-entropy and minimisation corresponds to maximisation of negative cross-entropy.
@@ A.2 Flow Matching as a Special Case of Variational Flow Matching
$=\mathbb{E}_{t,x,x_{1}}\left[\left\|u_{t}(x\mid x_{1})-v_{t}^{\theta}(x)\right\|_{2}^{2}\right]+,$ (35)
@@ A.3 Decomposition of the Flow
Applying this fact to eq. 38 we obtain:
@@ A.4 CatFlow
If $\theta_{t}(x)$ is permutation equivariant w.r.t $S_{|\mathcal{V}|}$, then so is $v_{t}$.
@@
$v_{t}(\pi\cdot x)=\frac{\theta_{t}(\pi\cdot x)-\pi\cdot x}{1-t}=\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t},$
@@
$\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t}=\pi\cdot\frac{(\theta_{t}(x)-x)}{1-t}=\pi\cdot v_{t}(x),$
@@
Let $p_{0}$ be an exchangeable distribution – e.g. a standard normal distribution – and $\theta_{t}(x)$ be permutation equivariant. Then, all permutations of graphs are generated with equal probability.
@@
[…]
@@ A.5 Derivation of the Score Function
[…]
@@ A.6 Stochastic Dynamics with Variational Flow Matching
we can rewrite eq. 50 in the following form:
@@
Importantly, the function $g_{t}$ does not affect the distribution path $p_{t}(x)$: it only changes the stochasticity of the trajectories. In the extreme case when $g_{t}\equiv 0$, the SDE in eq. 50 coincides with the ODE in eq. 49.
@@
The SDE in eq. 55 only depends on the score function, so in score-based models, the aim is to learn the score function.
Now, let us note that similarly to vector field $u_{t}(x)$, the drift term $\tilde{u}_{t}(x)$ in eq. 52 can be expressed in terms of an end point distribution $p_{t}(x_{1}\mid x)$. Consequently, having a variational approximation of end point distribution $q_{t}^{\theta}(x_{1}\mid x)$, allows us to construct an approximation of the drift term $\tilde{u}_{t}(x)$:
@@
Then, eq. 58 is not just simply a new dynamic, it is a family of dynamics that admits deterministic dynamics as a special case when $g_{t}\equiv 0$. Importantly, the only thing we need to construct the stochastic process in eq. 58 is an approximation of the end point distributions $q_{t}^{\theta}(x_{1}\mid x)$. Additionally we know that when $p_{t}(x_{1}\mid x)=q_{t}^{\theta}(x_{1}\mid x)$, $\tilde{u}_{t}(x)=\tilde{v}^{\theta}_{t}(x)$, the processes coincide for all functions $g_{t}$. Therefore, we can train the model with the same objective as in VFM.

### v2

@@ 4 Flow Matching and Score Matching: Bridging the Gap
VFM admits FM as a special case, under certain assumptions on $u_{t}(x\mid x_{1})$, when the variational approximation is Gaussian. Formally, the following holds (see theorem˜3 in section˜A.2 for the proof):
@@
Flow matching lipman2023flow is inspired by score-based models song2020score and shares strong connections with them. This leads us to two observations. The first is that it should be possible to define a variational parameterization of the score function in diffusion models that is analogous to the one in VFM. The second is that we can build on existing results that show that many diffusion model objectives, including the standard flow matching objective with a linear interpolant, can be expressed as special cases of a general weighted loss function kingma2021variational ; kingma2023understanding . Here we simililarly define a bound on the log-likelihood in terms of a reweighted VFM objective.
In score-based models, the objective is to approximate the score function $\nabla_{x}\log p_{t}(x)$ with a function $s^{\theta}_{t}(x)$. A connection to VFM becomes apparent by observing that the score function can also be expressed as an expectation with respect to $p_{t}(x_{1}\mid x)$ (see section˜A.5 for derivation):
@@
It is now clear that $s_{t}^{\theta}(x)=\nabla_{x}\log p_{t}(x)$ when $q_{t}^{\theta}(x_{1}\mid x)=p_{t}(x_{1}\mid x)$. This suggests that there exists a variational formulation of score-based models that is entirely analogous to VFM. Indeed, existing work on continuous diffusion for categorical data dieleman2022continuous defines a parameterization of the score function of this form (see section˜5 for a more detailed discussion).
Following song2020score ; anderson1982reverse , we can construct stochastic generative dynamics $dx=\tilde{v}_{t}^{\theta}(x)dt+g_{t}dw$ to approximate the true dynamics $dx=\tilde{u}_{t}(x)dt+g_{t}dw$ (see details in section˜A.6), with
@@
This connection has two important implications. First, it shows that learning a variational approximation $q^{\theta}_{t}(x_{1}\mid x)$ can be used to define both deterministic and stochastic dynamics, whereas flow matching typically considers deterministic dynamics only (as flows are viewed through the lens of ODEs). Second, it enables us to show that a reweighted version of the VFM objective provides a bound on the log-likelihood of the model. This result, inspired by kingma2023understanding , provides another theoretical motivation for learning using the VFM objective.
@@
We provide a proof in section˜A.7. We further note that that applying the same linearity condition that we discussed in section˜3.2 to the conditional score function maintains all the same connections with score-based models.
@@ 5 Related Work
Several approaches to diffusion models have been developed for graph generation. In vignac2022digress , the authors define a Markov process that progressively edits graphs by adding or removing edges and altering node or edge categories and is trained using a graph transformer network, that reverses this process to predict the original graph structure from its noisy version. This approach breaks down the complex task of graph distribution learning into simpler node and edge classification tasks. Moreover, jo2022scorebased proposes a score-based generative model for graph generation using a system of stochastic differential equations (SDEs). The model effectively captures the complex dependencies between graph nodes and edges by diffusing both node features and adjacency matrices through continuous-time processes. These are non-autoregressive graph generation approaches that perform on par with autoregressive ones, such as in liu2018constrained ; liao2019efficient ; mercado2021graph . Other non-autoregressive approaches worth mentioning are lippe2020categorical ; madhawa2019graphnvp ; luo2021graphdf .
There is also work on diffusion-based models for discrete data in the form of text and other sequential data austin2021structured ; luo2021graphdf ; yang2023diffsound ; inoue2023layoutdm . Though not explicitly formulated in terms of a variational perspective, the work on continuous diffusion for categorical data dieleman2022continuous arrives at a an approach that is closely related to that of CatFlow. This work defines a diffusion in the embedding space of a transformer-based language model. It defines an approximation of the score function as an expected value of a conditional score function as in eq.˜23, which leads to an expression for the learned score function in terms of a mean embedding, analogous to the one we obtain in eq.˜15. Where this approach differs from CatFlow, other than in that it defines a flow in the embedding space of language models, is in how the objective is defined. The authors also minimize a cross-entropy loss, but employ a time-warped objective, similar to the general weighted objective proposed in kingma2021variational , where the warping is optimized to ensure that the entropy of predictions decreases linearly when moving from noise to data in uniform time.
@@
Recently, two flow-based methods for discrete generative modeling have been proposed, which differ both in terms of technical approach and intended use case from the work that we present here. [^2] Flow matching and diffusion models have also been proposed for geometric graph generation, e.g. in klein2023equivariant ; song2023equivariant and hoogeboom2022equivariant ; trippe2022diffusion respectively, but since these approaches are continuous (as they generate coordinates based on some conformer) they consider a fundamentally different task than the ones we consider here.
In stark2024dirichlet , a Dirichlet flow framework for DNA sequence design is introduced, utilizing a transport problem defined over the probability simplex, similar to diffusion on simplices proposed in richemond2022categorical . This approach differs from CatFlow in that it represents the conditional probability path $p_{t}(x\mid x_{1})$ using a Dirichlet distribution. This implies that points $x$ are constrained to the simplex, which is not the case for CatFlow. Dirichlet Flows have not been evaluated on graph generation, but we did carry out preliminary experiments based on the released source code. We compare to this approach in our experiments.
In campbell2024generative , Discrete Flow Models (DFMs) are introduced. DFMs use Continuous-Time Markov Chains to enable flexible and dynamic sampling in multimodal generative modeling of both continuous and discrete data. Though sharing a goal, this approach differs significantly from CatFlow as in the end the resulting model does not learn a CNF, but rather generation through sequential sampling from a time-dependent categorical distribution. As in the case of Dirichlet flows, no evaluation on graph generation was performed.
The switch to the variational perspective is inspired by vignac2022digress , showing significant improvement through viewing the dynamics as a classification task over end points. However, CatFlow is still a continuous model, and integrates – rather than iteratively samples – during generation.
@@
GraphVAE simonovsky2018graphvae 0.130 0.170 0.050 0.350 0.980 0.540
GNF liu2019graph 0.030 0.100 0.001 0.200 0.200 0.110
EDP-GNN niu2020permutation 0.052 0.093 0.007 0.053 0.144 0.026
GDSS jo2022scorebased 0.021 0.024 0.007 0.045 0.086 0.007
@@ 6 Experiments
We evaluate CatFlow in three sets of experiments. First, we consider an abstract graph generation task proposed in martinkus2022spectre , where the goal of this task is to evaluate if CatFlow is able to capture the topological properties of graphs. Second, we consider two common molecular benchmarks, QM9 ramakrishnan2014quantum and ZINC250k irwin2012zinc , consisting of small and (relatively) large molecules respectively. This task is chosen to see if CatFlow can learn semantic information in graph generation, such as molecular properties. Finally, we perform an ablation comparing CatFlow to standard flow matching, specifically in terms of generalization. The experimental setup and model choices are provided in appendix˜D.
@@
MoFlow zang2020moflow 91.36 98.65   4.467 63.11 99.99 20.931
EDP-GNN niu2020permutation 47.52 99.25   2.680 82.97 99.79 16.737
GraphEBM liu2021graphebm   8.22 97.90   6.143   5.29 98.79 35.471
GDSS jo2022scorebased 95.72 98.46   2.900 97.01 99.64 14.656
Digress vignac2022digress 99.00 96.20 - - - -
Flow Matching lipman2023flow 94.10 98.20   5.155 94.01 96.68 18.764
Dirichlet FM stark2024dirichlet 99.10 98.15   0.888 97.52 99.20 14.222
@@ 6.1 Abstract graph generation
We first evaluate CatFlow on an abstract graph generation task, including synthetic and real-world graphs. We consider 1) Ego-small (200 graphs), consisting of small ego graphs drawn from a larger Citeseer network dataset sen2008collective , 2) Community-small (100 graphs), consisting of randomly generated community graphs, 3) Enzymes (587 graphs), consisting of protein graphs representing tertiary structures of the enzymes from schomburg2004brenda , and 4) Grid (100 graphs), consisting of 2D grid graphs. We follow the standard experimental setup popularized by you2018graphrnn and hence report the maximum mean discrepancy (MMD) to compare the distributions of degree, clustering coefficient, and the number of occurrences of orbits with 4 nodes between generated graphs and a test set. Following jo2022scorebased , we also use the Gaussian Earth Mover’s Distance kernel to compute the MMDs instead of the total variation.
The results of the Ego-small and Community-small tasks are summarized in table˜1, and additional results (and error bars) are provided in appendix˜C. The results indicate that CatFlow is able to capture topological properties of graphs, and performs well on abstract graph generation.
@@ 6.2 Molecular Generation: QM9 & ZINC250k
Molecular generation entails designing novel molecules with specific properties, a complex task hindered by the vast chemical space and long-range dependencies in molecular structures. We evaluate CatFlow on two popular molecular generation benchmarks: QM9 and ZINC250k ramakrishnan2014quantum ; irwin2012zinc .
We follow the standard setup – e.g. as in shi2020graphaf ; luo2021graphdf ; vignac2022digress ; jo2022scorebased – of kekulizing the molecules using RDKit Landrum2016RDKit2016_09_4 and removing the hydrogen atoms. We sample 10,000 molecules and evaluate them on validity, uniqueness, and Fréchet ChemNet Distance (FCD) – evaluating the distance between data and generated molecules using the activations of ChemNet preuer2018frechet . Here, validity is computed without valency correction or edge resampling, hence following zang2020moflow rather than shi2020graphaf ; luo2021graphdf , as is more reasonable due to the existence of formal charges in the data itself. We do not report novelty for QM9 and ZINC250k, as QM9 is an exhaustive list of all small molecules under some chemical constraint and all models obtain (close to) 100% novelty on ZINC250k. [^3] CatfFlow obtains 49% novelty on QM9.
The results are summarized in table˜2 and samples from the model are shown in fig.˜2. CatFlow obtains state-of-the-art results on both QM9 and ZINC250k, virtually obtaining perfect performance on both datasets. It is worth noting that CatFlow also converges faster than flow matching and is not computationally more expensive than any of the baselines either during training or generation.
@@ 6.3 CatFlow Ablations
Figure 3: Ablation results. We compare standard flow matching and CatFlow. We visualize performance degradation in terms of a score, which is the percentage of molecules that is valid and unique, for a varying number of layers and percentage of the training data.
@@
In fig.˜3 we report a score, which is the percentage of generated molecules that is valid and unique. CatFlow not only outperforms regular flow matching in the large model and full data setting, but is also significantly more robust to a decrease in model-size and data. Moreover, we observe significantly faster convergence (curves not shown). We hypothesize this is a consequence of the optimization procedure not exploring ‘irrelevant’ paths that do not point towards the probability simplex.
@@ A.1 Derivation of the Variational Flow Matching Objective
We derive eq.˜9 that states the equivalence of the optimization of the VFM objective and minimization of the KL divergence between the true endpoint distribution $p_{t}(x_{1}\mid x)$ and the variational approximation $q^{\theta}_{t}(x_{1}\mid x)$. Note that
@@
Therefore, the second part of eq.˜27 corresponds to negative cross-entropy and minimisation corresponds to maximisation of negative cross-entropy.
@@ A.2 Flow Matching as a Special Case of Variational Flow Matching
$=\mathbb{E}_{t,x,x_{1}}\left[\left\|u_{t}(x\mid x_{1})-v_{t}^{\theta}(x)\right\|_{2}^{2}\right]+C,$ (35)
@@ A.3 Decomposition of the Flow
Applying this fact to eq.˜38 we obtain:
@@ A.4 CatFlow
If $\mu_{t}(x)$ is permutation equivariant w.r.t $S_{|\mathcal{V}|}$, then so is $v_{t}$.
@@
$v_{t}(\pi\cdot x)=\frac{\mu_{t}(\pi\cdot x)-\pi\cdot x}{1-t}=\frac{\pi\cdot\mu_{t}(x)-\pi\cdot x}{1-t},$
@@
$\frac{\pi\cdot\mu_{t}(x)-\pi\cdot x}{1-t}=\pi\cdot\frac{(\mu_{t}(x)-x)}{1-t}=\pi\cdot v_{t}(x),$
@@
Let $p_{0}$ be an exchangeable distribution – e.g. a standard normal distribution – and $\mu_{t}(x)$ be permutation equivariant. Then, all permutations of graphs are generated with equal probability.
@@
By result 1, we know that $\mu_{t}(x)$ being permutation equivariant implies that $v_{t}(x)$ is permutation equivariant. Moreover, if we let $\Gamma(x):=x+\int_{0}^{1}v_{t}(x_{t})dt$, the for all $\pi\in S_{n}$ we have that
$\Gamma(\pi\cdot x)=\pi\cdot x+\int_{0}^{1}v_{t}(\pi\cdot x_{t})dt=\pi\cdot x+\int_{0}^{1}\pi\cdot v_{t}(x_{t})dt=\pi\cdot\Gamma(x),$
@@ A.5 Derivation of the Score Function
$=\frac{1}{p_{t}(x)}\nabla_{x}\int p_{t}(x\mid x_{1})p(x_{1})dx_{1}$ (44)
$=\frac{1}{p_{t}(x)}\int p(x_{1})\nabla_{x}p_{t}(x\mid x_{1})dx_{1}$ (45)
$=\int\frac{p_{t}(x\mid x_{1})p(x_{1})}{p_{t}(x)}\nabla_{x}\log p_{t}(x\mid x_{1})dx_{1}$ (46)
$=\int p_{t}(x_{1}\mid x)\nabla_{x}\log p_{t}(x\mid x_{1})dx_{1}$ (47)
@@ A.6 Stochastic Dynamics with Variational Flow Matching
we can rewrite eq.˜50 in the following form:
@@
Importantly, the function $g_{t}$ does not affect the distribution path $p_{t}(x)$: it only changes the stochasticity of the trajectories. In the extreme case when $g_{t}\equiv 0$, the SDE in eq.˜50 coincides with the ODE in eq.˜49.
@@
The SDE in eq.˜55 only depends on the score function, so in score-based models, the aim is to learn the score function.
Now, let us note that similarly to vector field $u_{t}(x)$, the drift term $\tilde{u}_{t}(x)$ in eq.˜52 can be expressed in terms of an end point distribution $p_{t}(x_{1}\mid x)$. Consequently, having a variational approximation of end point distribution $q_{t}^{\theta}(x_{1}\mid x)$, allows us to construct an approximation of the drift term $\tilde{u}_{t}(x)$:
@@
Then, eq.˜58 is not just simply a new dynamic, it is a family of dynamics that admits deterministic dynamics as a special case when $g_{t}\equiv 0$. Importantly, the only thing we need to construct the stochastic process in eq.˜58 is an approximation of the end point distributions $q_{t}^{\theta}(x_{1}\mid x)$. Additionally we know that when $p_{t}(x_{1}\mid x)=q_{t}^{\theta}(x_{1}\mid x)$, $\tilde{u}_{t}(x)=\tilde{v}^{\theta}_{t}(x)$, the processes coincide for all functions $g_{t}$. Therefore, we can train the model with the same objective as in VFM.